OpenStudy (anonymous):

two cartons of milk and one sandwich cost 3$ and one carton of milk and two sandwiches cost 4.20$. whats the cost of one carton of milk

OpenStudy (anonymous):

2m+s=3 m+2s=4.2 Then you would solve using substitution or combining the equations, etc

OpenStudy (anonymous):

my calculator has an option that solves two line equations.

OpenStudy (anonymous):

Using substitution you get: s = 3-2m m + 2(3-2m) = 4.2 m+6-4m=4.2 -3m=-1.8 m=0.60 had a typo earlier